OpenStudy (anonymous):

Evaluate lim (cosh^-1(x)-ln(x)) as x goes to infinity

OpenStudy (dan815):

arccosh or 1/cosh(x)?

OpenStudy (anonymous):

arccosh

OpenStudy (dan815):

so remember cosh can be written in terms of e and ln is the inverse of e

OpenStudy (anonymous):

I don't remember can you write to me ? please

ganeshie8 (ganeshie8):

Let \(y=\arccos hx \) \(\implies x = \cosh y = \dfrac{e^y + e^{-y}}{2}\) \(\implies 2x = e^y+e^{-y}\) \(\implies e^{2y} -2xe^y +1=0\) this is a quadratic, solving should give you \(y=\arccos h x = \ln(x + \sqrt{x^2-1)}\)

ganeshie8 (ganeshie8):

plug that thing in the given expression rest should be easy

OpenStudy (anonymous):

thank you so much.. it is helpful

ganeshie8 (ganeshie8):

np, here is the rest : \(\lim\limits_{x\to\infty} \cosh^{-1}(x)-\ln(x) \) \(=\lim\limits_{x\to\infty} \ln(x+\sqrt{x^2-1})-\ln(x) \) \(=\lim\limits_{x\to\infty} \ln(\dfrac{x+\sqrt{x^2-1}}{x}) \) \(=\lim\limits_{x\to\infty} \ln(1+\sqrt{1-\dfrac{1}{x^2}}) \) \(= \ln(\lim\limits_{x\to\infty} 1+\sqrt{1-\dfrac{1}{x^2}}) \) \(=\ln(1+1)\)

OpenStudy (anonymous):

Thank you soooooooooooo much

OpenStudy (anonymous):

ln(1+1)